OpenStudy (anonymous):

Natalie and her sister opened savings accounts at the same time. Natalie deposited $100 and will deposit $10 each month. Her sister deposited $25 and will deposit $25 each month. When will Natalie have less money in her account than her sister?

OpenStudy (anonymous):

natilie: y=100+10x her sister: y=25+25x